OpenStudy (anonymous):

help please @ganeshie8 @mathmale @whpalmer4 @AravindG @helpme1.2

OpenStudy (anonymous):

hold on

OpenStudy (mathmale):

Look carefully at the given graph. What is the slope of the line shown in your diagram? What is the y-intercept?

OpenStudy (anonymous):

8?

OpenStudy (mathmale):

I'm sure you've seen the equation used to calculate the slope of a straight line. You are given a graph that clearly shows 2 points on that line. Use the slope formula to find the slope of this line. The slope is not 8.

OpenStudy (anonymous):

ooo ok

OpenStudy (mathmale):

Please find the slope, using the slope formula. Look up this formula if you don't remember it.

OpenStudy (anonymous):

2 of 3/6

OpenStudy (mathmale):

And so the slope of this straight line is m = ??

OpenStudy (mathmale):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

thats what i got

OpenStudy (mathmale):

the slope is actually 1/2, not 2. Would you mind showing me your calculations, so that i could give you feedback on them if need be? m=?

OpenStudy (anonymous):

\[\frac{ 7-4 }{ 8-2 } =\frac{ 3 }{ 6 } or 2 \]

OpenStudy (anonymous):

huh?

OpenStudy (mathmale):

Excuse me. In your calculation you write 3/6, and then you write "or 2." Mind explaining that?

OpenStudy (anonymous):

i simplified

OpenStudy (mathmale):

Do you really mean to say that 3/6 = 2? Think about that.

OpenStudy (mathmale):

Maya, 3/6 reduces to 1/2, not to 2. Therefore the slope of this straight line is m = 1/2. That eliminates which of the four answer choices? Now look at the two points on the line identified by their coordinates. Substitute the appropriate coordinates and your slope into the point-slope formula for the equation of a straight line. Which of the four answer choices could be the one that describes this straight line? There might be more than one correct answer.

OpenStudy (anonymous):

d,e.f

OpenStudy (whpalmer4):

The point-slope formula is used to write the equation for a line passing through a given point \((x_1,y_1)\) with slope \(m\). The formula is\[y-y_1 = m(x-x_1)\] We know that the slope here is \(m = \frac{1}{2}\), so the equation will look like \[y-y_1 = \frac{1}{2}(x-x_1)\] Answers D, E, and F can all be eliminated as incorrect because they feature a slope of 2 rather than 1/2. You know two points on the line: (2,4) and (8,7). Construct a line equation with slope m = 1/2 using the point-slope formula for both of those points. You should find them in the remaining answers if you do the work correctly.